Here is the problem I am trying to solve: The square of a positive number is six more than five times the positive number. Find the number.

Answer by mananth(16949)

let the number be x
x^2 =5x+6
x^2-5x-6=0
x^2-6x+x-6=0
x(x-6)+1(x-6)=0
(x-6)(x+1)
x=6 OR -1